Part 1 - A payday loan company charges 2.2 percent interest for a two-week period. What would be the annual interest rate from that company? (Assume an even 52 weeks per year. Enter your answer as a percent rounded to 1 decimal place.)

Part 2 - A worker's primary goal should be to

Multiple Choice

  • Pay his or her taxes using estimates for income and deductions.

  • Pay no income taxes.

  • Pay the average tax rate for people working in his or her industry.

  • Pay his or her fair share of taxes while taking advantage of appropriate tax benefits.

  • Pay no taxes of any type.
Answer #1

Part 1

Interest rate for 2 Week = 2.2%

Number of weeks in a year = 52

So the Interest rate for the Year will be= 2.2*52 weeks/2 weeks = 57.2%

Example:

If the loan amount is 1000

2 Weeks interest will be : 1000*2.2% = 22

So the Interest amount for the year = 22*52/2 = 572

ie Rate of Interest will be = 572*100/1000 = 57.2%

Part 2

Primary goal of the worker will be: Pay no taxes of any type.

Common people are not willing to pay taxes at all, as its a deduction from their income. Person with normal prudence will be interested to spend less money than what they earn, they try to reduces the expenses by reducing the unnecessary expenses. As the Taxes are payment to the government and they are not get directly benefited from the same, peoples are normally unwilling to pay taxes of any types unless it is compulsory.